OpenStudy (anonymous):

If x ≠ 0 and x ≠ 14, what is the solution to the equation below. 2/x - 14 = 3 / 4x **Please show how you got your answer.

OpenStudy (anonymous):

\[\frac{ 2 }{ x - 14 } = \frac{ 3 }{ 4x }\]

OpenStudy (anonymous):

move everything to one side, find a common denominator, then solve for x